# Solar Cruiser Disturbance Torque Estimation and Predictive Momentum Management

### v1

@@ 5.3 Proposed MPC-based Momentum Management Supported by KF Disturbance Estimate
$\mbf{Q}_{N}$ $\mbf{1}_{13\times 13}$
$\mbf{R}$ $\text{diag}(1,1,10^{6})$
$\tilde{\mbf{R}}$ $100\cdot\mbf{1}_{2\times 2}$
@@
It is worth noting that the MPC evaluates RCD inputs as continuous values between $u^{\text{RCD}}_{\text{min}}$ and $u^{\text{RCD}}_{\text{max}}$, but the actual applied input is quantized into the full on/off value with pulse length $t_{c}$ using PWM quantization as in Eq. (8).
To demonstrate the importance of disturbance knowledge in the MPC framework, simulation results with and without the Kalman filter disturbance estimate knowledge in MPC are shown in Fig. 4, where no threshold is used ($\beta^{\text{AMT}}_{\text{thresh}}=\beta^{\text{RCD}}_{\text{thresh}}=0$). The result in blue labeled “nominal MPC” uses the nominal MPC implementation without disturbance knowledge, where the MPC prediction model uses $\mbf{w}_{j|t_{k}}=\mbf{0}$, for $j=0,1,\ldots,N-1$. The result in red labeled “KFdist MPC” includes the Kalman filter estimate disturbance within the MPC prediction model, where $\mbf{w}_{j|t_{k}}=\hat{\mbf{w}}^{+}_{k}$, for $j=0,1,\ldots,N-1$. Although both of the MPC policies perform successful momentum management under a $6^{\circ}$ slew, the nominal MPC results in the angular momentum of two RWs stabilizing near their saturation limits. Conversely, the MPC implementation incorporating the disturbance estimate exhibits a significant performance improvement, driving all RW angular momentum down to values safely within the specified soft constraint boundaries, thereby reserving greater control authority.
@@
Figure 4: Simulation results using the proposed MPC momentum management strategy under a 66 deg slew with and without (nominal) the disturbance estimate knowledge in prediction model. The black dashed lines in (b) denote the $25\%$ soft constraint on 4-RWs angular momentum.
An additional actuation threshold can be applied on the MPC inputs to filter out minor actuation with minimal loss in momentum management performance. Leveraging the recursive nature of the MPC, an input activation threshold is applied to trim out minor actuation demanded by MPC, and further improve actuator efficiency and mitigate noise. A set of results are presented in Fig. 5, demonstrating the design choice of actuation thresholds. The result in blue (labeled “MPC-thrA1R3”) uses a $10\%$ AMT threshold ($\beta^{\text{AMT}}_{\text{thresh}}=0.1$) and $30\%$ RCD threshold ($\beta^{\text{RCD}}_{\text{thresh}}=0.3$), which means that when MPC demands an AMT input less than $10\%$ of the distance the AMT can move in one direction in one time step ($10\%$ of $0.05$ m), the AMT is held at its current position for the next time step, and the RCD input is set to zero when the MPC-demanded input is less than $30\%$ of the RCD “on” torque value. The result in red (labeled “MPC-thrA2R6”) uses a $20\%$ AMT threshold ($\beta^{\text{AMT}}_{\text{thresh}}=0.2$) and $60\%$ RCD threshold ($\beta^{\text{RCD}}_{\text{thresh}}=0.6$) on the MPC-demanded inputs. Figures 5(a) and 5(b) show that the design choice of the applied thresholds do not degrade momentum management performance, which is further illustrated in the plot of the control inputs in Fig. 5(c) and the zoomed in control input plot of Fig. 5(d). A comparison of actuation usage among the the three MPC policies with different actuation threshold is included in Table 4. The performance metric of control actuation effort is evaluated by the number of RCD on-off cycles, the total time the RCDs are turned “on”, the total AMT travel distance in each translation axis, and the sum of the total AMT travel distance across both axes.
@@
Figure 5: Threshold tuning using $10\%$ AMT threshold and $30\%$ RCD threshold (blue), and using $20\%$ AMT threshold and $60\%$ RCD threshold (red). The zoomed-in plot in (d) demonstrates the first $2000$ seconds of slew maneuver and the PWM-quantized RCD actuation pulsing at every time step.
@@
AMT/RCD Threshold AMT 0% / RCD 0% AMT 10% / RCD 30% AMT 20% / RCD 60%
RCD Cycle (#\#) 74 69 48
RCD On Time (sec) 3850 3846 3593
AMT Dist 1 (cm) 33.85 24.21 26.33
AMT Dist 2 (cm) 100.11 95.50 93.43
Sum of AMT Dist (cm) 133.96 119.71 119.76
Figure 6 illustrates the disturbance torque estimates generated by the Kalman filter for the three MPC test cases. The black dashed lines are the true disturbances, accounting for the torque generated by the SRP due to the offset between the CM and CP. While the exact magnitude of the estimated disturbance torque does not exactly match the true disturbance torque, the estimate is reasonably accurate, and clearly assists with the MPC-based momentum management strategy, as shown in Fig. 4. It is worth noting that the disturbance torque estimate generated by the Kalman filter will account for all model inaccuracies in practice (e.g., nonlinearities, discretization approximations), which could explain the difference between the estimated and true disturbance torque.
Figure 6: Kalman filter disturbance estimate values used in the MPC momentum management with $10\%$ AMT and $30\%$ RCD threshold (blue), $20\%$ AMT and $60\%$ RCD threshold (red), and no threshold (yellow).
@@ 5.4 State-of-the-Art Comparison
A $3^{\circ}$ slew simulation is executed using the MPC framework to directly compare its actuation efficiency to that of NASA’s state-of-the-art method (Inness et al., 2023; Tyler et al., 2023). For this simulation, MPC uses the same tuning parameters as the simulations in Fig. 5 with $20\%$ AMT threshold and $60\%$ RCD threshold.
Figure 7 includes the comparison of simulation results using NASA’s method (Inness et al., 2023; Tyler et al., 2023) and the proposed MPC approach with thresholds. In Fig. 7(d), the MPC proactively actuates the AMT and RCDs to avoid angular momentum growth, as shown in Figures 7(b) and 7(c). A quantitative comparison of the control actuation usage is included in Table 5. The proposed MPC policy achieves a significant reduction in the total AMT travel distance. In contrast, the total RCD activation time is comparable between both methods. The PWM-quantization evenly distributes the input across every time step, as opposed to the longer singular “on” pulse with a long “off” period when using NASA’s benchmark method. Although the MPC results in higher RCD on-off cycles due to this inherent PWM quantization, dividing a long activation command into multiple short pulses is not inherently detrimental, as it mitigates the risk of potentially overheating the actuator associated with excessively long RCD “on” commands. Future work could investigate the design of an actuation mechanism capable of grouping these short MPC-generated pulses into a single, longer RCD activation event according to the mission requirements and hardware limitations.
@@
Figure 7: Comparison of simulation results using the proposed MPC momentum management strategy versus NASA’s state-of-the-art method under a $3^{\circ}$ slew.
Table 5: Control effort usage with NASA’s state-of-the-art momentum management method and the proposed MPC method with $20\%$ AMT and $60\%$ RCD threshold under a slew from ${\boldsymbol{\theta}}_{0}=\big[-3^{\circ}\,\,\,0\,\,\,0\big]^{\mathsf{T}}$ to ${\boldsymbol{\theta}}_{d}=\mbf{0}$.
Controller NASA Solar Cruiser MPC w/ Threshold
RCD Cycle (#\#) 1 49
RCD On Time (sec) 3300 3538
AMT Dist 1 (cm) 30.64 20.40
AMT Dist 2 (cm) 166.56 72.58
Sum of AMT Dist (cm) 197.20 92.98
@@ 6 Conclusions
This paper presented a novel Kalman filter augmented MPC framework specifically designed for the challenging momentum management task of NASA’s Solar Cruiser. The integrated estimation framework proposed in this work plays a crucial role, providing real-time state and disturbance estimates that not only characterize the external disturbance torque but also capture the dynamic discrepancies between the linear prediction model and the highly nonlinear spacecraft system. This estimate closes the modeling gap needed to enable model predictive control for this momentum management application. Building upon a previously-developed MPC architecture (Shen and Caverly, 2025b), the policy was rigorously formulated to be computationally feasible, utilizing off-the-shelf QP solvers to ensure real-time implementation capability within the limited onboard hardware.
Simulation results demonstrated the proposed MPC-based momentum management policy’s superior performance and robustness. The disturbance estimate was shown to be essential in achieving reliable MPC prediction and bounded momentum management. Furthermore, the proposed MPC policy successfully managed angular momentum growth under maneuvers that exceed the capability of NASA’s state-of-the-art method designed for Solar Cruiser, establishing a larger operational slew envelope. The framework also proved its efficiency by demonstrating reduced actuator usage through a lower AMT travel distance and optimized RCD usage compared to the benchmark method. This improvement has the potential to enable greater solar sail mission longevity.
Future work on this topic could be the investigation of tracking larger, more dynamic slew maneuvers and the development of methodologies to proactively unload angular momentum in preparation for upcoming high-demand maneuvers. Additional work towards the implementation of the proposed method on flight hardware and software will also be pursued to move towards its implementation on future solar sail missions.
@@ Acknowledgments
This material is based upon work supported by NASA under award No. 80NSSC25M7060, as well as a study grant from Chung Cheng Institute of Technology, National Defense University, Taiwan (R.O.C.).

### v2

@@ 5.3 Proposed MPC-based Momentum Management Supported by KF Disturbance Estimate
$\mbf{Q}_{N}$ $10\cdot\mbf{Q}$
$\mbf{R}$ $\text{diag}(1,1,5\times 10^{6})$
$\tilde{\mbf{R}}$ $2000\cdot\mbf{1}_{2\times 2}$
@@
It is worth noting that the MPC evaluates RCD inputs as continuous values between $\pm\tau^{\text{RCD}}_{b3,\text{on}}$, but the actual applied input is quantized into the full on/off value with pulse length $t_{c}$ using PWM quantization as in Eq. (11). The current SRP force $\mbf{f}_{b}^{\text{SRP}}(t_{k})$ and RCD on torque $\tau^{\text{RCD}}_{b3,\text{on}}(t_{k})$ used in MPC are updated at every momentum management timestep $t_{k}$, while the Kalman filter provides the state and disturbance estimates. MPC’s assumption that these values are constant across the prediction horizon further shows the robustness of the proposed method. The relaxation of the on-off RCD actuation constraints allows for the use of off-the-shelf QP solvers that can solve the optimization problem efficiently. The MATLAB function quadprog with its default settings is used. The mean QP solution time across $600$ momentum management timesteps is $32.86$ ms, and the numerical integration time for the $N=10$ LTV prediction model is $39.80$ ms, representing the average of five simulation sets of ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,15^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ maneuver. For reference, these computations are performed on a desktop computer with a 13th Gen Intel Core i5-13400 @ 2.5 GHz with 24 GB memory and the code is run in Matlab 2024b.
To demonstrate the importance of disturbance knowledge in the MPC framework, simulation results with and without the Kalman filter disturbance estimate knowledge in MPC are shown in Fig. 4, where no threshold is used ($\beta^{\text{AMT}}_{\text{thresh}}=\beta^{\text{RCD}}_{\text{thresh}}=0$). The result in blue labeled “nominalMPC” uses the nominal MPC implementation without disturbance knowledge, where the MPC prediction model uses $\mbf{w}_{j|t_{k}}=\mbf{0}$, for $j=0,1,\ldots,N-1$. The result in red labeled “KFMPC” includes the Kalman filter estimate disturbance within the MPC prediction model, where $\mbf{w}_{j|t_{k}}=\hat{\mbf{w}}^{+}_{k}$, for $j=0,1,\ldots,N-1$. Although both of the MPC policies perform successful momentum management under an attitude hold at ${\boldsymbol{\theta}}_{0}={{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$, the nominal MPC results in the angular momentum of two RWs stabilizing near their saturation limits. Conversely, the MPC implementation incorporating the disturbance estimate exhibits a significant performance improvement, driving all RW angular momentum down to values safely within the specified soft constraint boundaries, thereby reserving greater control authority. An attitude hold at ${\boldsymbol{\theta}}=\mbf{0}$ (which has a higher disturbance torque) and other slew maneuvers have been tested without the disturbance estimate, all of which resulted in RW saturation and instability. This further shows that the disturbance estimate is critical to the performance of MPC-based momentum management.
@@
Figure 4: Simulation results using the proposed MPC momentum management strategy under an attitude hold of ${\boldsymbol{\theta}}_{0}={{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ with and without (nominal) the disturbance estimate knowledge in prediction model. The black dashed lines in (b) denote the $25\%$ soft constraint on 4-RWs angular momentum.
The proposed MPC policy with Kalman filter estimation framework is used to perform the maneuver sequence tracking initial attitude of ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,15^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$, following the desired hold-slew-hold-slew-hold trajectory. An additional actuation threshold can be applied on the MPC inputs to filter out minor actuation with minimal loss in momentum management performance. Leveraging the recursive nature of the MPC, an input activation threshold is applied to trim out minor actuation demanded by MPC, and further improve actuator efficiency and mitigate noise. A set of results are presented in Fig. 5, demonstrating the design choice of actuation thresholds. The result in blue (labeled “MPC-thrA2R6”) uses a $20\%$ AMT threshold ($\beta^{\text{AMT}}_{\text{thresh}}=0.2$) and $60\%$ RCD threshold ($\beta^{\text{RCD}}_{\text{thresh}}=0.6$), which means that when MPC demands an AMT input less than $20\%$ of the distance the AMT can move in one direction in one time step ($20\%$ of $0.05$ m), the AMT is held at its current position for the next time step, and the RCD input is set to zero when the MPC-demanded input is less than $60\%$ of the RCD “on” torque value. The result in red (labeled “MPC-thrA3R9”) uses a $30\%$ AMT threshold ($\beta^{\text{AMT}}_{\text{thresh}}=0.3$) and $90\%$ RCD threshold ($\beta^{\text{RCD}}_{\text{thresh}}=0.9$) on the MPC-demanded inputs. Figures 5(a) and 5(b) show that the design choice of the applied thresholds do not degrade momentum management performance, which is further illustrated in the plot of the control inputs in Fig. 5(c) and the zoomed in control input plot of Fig. 5(d). A comparison of actuation usage among the the three MPC policies with different actuation threshold performing the $15^{\circ}$ slew maneuver sequence over $60000$ seconds is included in Table 4. The performance metric of control actuation effort is evaluated by the number of RCD on-off cycles, the total time the RCDs are turned “on”, the total AMT travel distance in each translation axis, and the sum of the total AMT travel distance across both axes. The design choice of thresholds $\beta^{\text{AMT}}_{\text{thresh}}$ and $\beta^{\text{RCD}}_{\text{thresh}}$ can be determined by the operational characteristics and the expected lifetime of the actuators.
@@
Figure 5: Threshold tuning of the proposed MPC framework under a maneuver sequence tracking initial attitude ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,15^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ and return, using $20\%$ AMT threshold and $60\%$ RCD threshold (blue), and using $30\%$ AMT threshold and $90\%$ RCD threshold (red). The zoomed-in plot in (d) demonstrates the control inputs between $9000$ to $13000$ seconds of forward slew maneuver and the PWM-quantized RCD actuation pulsing at every time step.
@@
AMT/RCD Threshold AMT 10% / RCD 30% AMT 20% / RCD 60% AMT 30% / RCD 90%
RCD Cycle (#\#) 42 28 12
RCD On Time (sec) 6316 5554 4989
AMT Dist 1 (cm) 220.52 235.25 241.13
AMT Dist 2 (cm) 152.24 147.82 144.74
Sum of AMT Dist (cm) 372.76 383.07 385.87
Figure 6 illustrates the disturbance torque estimates generated by the Kalman filter for the three MPC test cases. The black dashed lines are the true disturbances, accounting for the torque generated by the SRP due to the non-ideal sail shape and the solar sail’s attitude. The forward slew and return slew are initiated at $10000$ and $20000$ seconds respectively, which results in the change of SRP disturbance torque. While the exact magnitude of the estimated disturbance torque does not exactly match the true disturbance torque, the estimate is reasonably accurate, and clearly assists with the MPC-based momentum management strategy, as shown in Fig. 4. It is worth noting that the disturbance torque estimate generated by the Kalman filter will account for all model inaccuracies in practice (e.g., nonlinearities, discretization approximations), which could explain the difference between the estimated and true disturbance torque. The disturbance torque in roll axis has a significantly smaller magnitude than the pitch/yaw axes and the other state estimates, making it difficult to observe and sensitive to measurement noise. However, the roll disturbance estimate still converges within the neighborhood of the true value, and allows MPC to adjust accordingly. Future work will investigate improving the observability of the roll disturbance estimate, through the use of more accurate measurements or the introduction of additional measurements.
Given the prediction model being linearized about the nominal slew trajectory, the MPC policies proactively take momentum management actuation once the slew maneuver arises in the prediction horizon (1000 seconds ahead in this case). Within the MPC formulation, the RCD torque magnitude, SRP force, estimated disturbance torque are assumed constant, with the AMT actuation modeled as a ZOH. In contrast, the simulation incorporates the nonlinear attitude dynamics, AMT motion dynamics, and the attitude-dependent nature of the SRP force, torque, and RCD effects. Despite the simplifications in the prediction model of MPC (which improves real-time feasibility), the recursive nature allows the controller to compensate for these discrepancies at every time step, demonstrating significant robustness against model uncertainties.
Figure 6: Kalman filter disturbance estimate values used in the MPC momentum management with $10\%$ AMT and $30\%$ RCD threshold (blue), $20\%$ AMT and $60\%$ RCD threshold (red), and $30\%$ AMT and $90\%$ RCD threshold (yellow). The black dashed line indicates the true SRP disturbance torque.
@@ 5.4 State-of-the-Art Comparison
A maneuver sequence tracking initial attitude of ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ following the desired hold-slew-hold-slew-hold trajectory is executed using the MPC framework to directly compare its actuation efficiency to that of NASA’s state-of-the-art method (Inness et al., 2023; Tyler et al., 2023). This threshold-based control policy is a recreation of the momentum management logic for Solar Cruiser based on the information provided by Inness et al. (2023); Tyler et al. (2023). While it serves as a functional approximation for the purposes of this study, it is not an exact replica the proprietary controller implemented on NASA’s flight hardware. For this simulation, MPC uses the same tuning parameters as the simulations in Fig. 5 with $30\%$ AMT threshold and $90\%$ RCD threshold.
Figure 7 includes the comparison of simulation results using NASA’s method (Inness et al., 2023; Tyler et al., 2023) and the proposed MPC approach with thresholds. In Fig. 7(d), the MPC proactively actuates the AMT and RCDs to avoid angular momentum growth, as shown in Figures 7(b) and 7(c). A quantitative comparison of the control actuation usage is included in Table 5. The proposed MPC policy achieves a significant reduction in AMT and RCD usage during the attitude hold. In contrast, a higher actuation usage during the slew maneuver provides a significantly more effective momentum management and a broader operation region (i.e., a larger range of slew maneuvers in which momentum management can be effectively performed. The PWM-quantization evenly distributes the input across every time step, as opposed to the longer singular “on” pulse with a long “off” period when using NASA’s benchmark method. Although the MPC results in higher RCD on-off cycles due to this inherent PWM quantization, dividing a long activation command into multiple short pulses is not inherently detrimental, as it mitigates the risk of potentially overheating the actuator associated with excessively long RCD “on” commands. Future work could investigate the design of an actuation mechanism capable of grouping these short MPC-generated pulses into a single, longer RCD activation event according to the mission requirements and hardware limitations.
@@
Figure 7: Comparison of simulation results using the proposed MPC momentum management strategy versus NASA’s state-of-the-art method under a maneuver sequence tracking initial attitude ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ and return.
Table 5: Control actuation usage with NASA Solar Cruiser’s state-of-the-art momentum management method and the proposed MPC method with $30\%$ AMT and $90\%$ RCD threshold under a maneuver sequence from ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ and return.
Time Slew Maneuver (0-30000 s) Attitude Hold (30000-60000 s)
Controller Solar Cruiser MPC w/ Threshold Solar Cruiser MPC w/ Threshold
RCD Cycle (#\#) 5 10 1 0
RCD On Time (sec) 3900 4695 1700 0
AMT Dist 1 (cm) 171.48 208.75 63.26 9.53
AMT Dist 2 (cm) 62.24 142.22 62.05 12.65
Sum of AMT Dist (cm) 233.72 350.97 125.31 22.18
@@ 6 Conclusions
This paper presented a novel Kalman filter augmented MPC framework specifically designed for the challenging momentum management task of NASA’s Solar Cruiser. The integrated estimation framework proposed in this work plays a crucial role, providing real-time state and disturbance estimates that not only characterize the external disturbance torque but also capture the dynamic discrepancies between the linear prediction model and the highly nonlinear spacecraft system. This estimate closes the modeling gap needed to enable model predictive control for this momentum management application. Building upon a previously-developed MPC architecture (Shen and Caverly, 2026), the policy was rigorously formulated to be computationally feasible, utilizing off-the-shelf QP solvers to ensure real-time implementation capability within the limited onboard hardware. Validation with a more realistic SRP model, while performing slew maneuver following and incorporating Solar Cruiser’s 4-RW configuration in this paper presents a key contribution towards the development of a practical implementation of the proposed MPC-based momentum management policy.
Simulation results demonstrated the proposed MPC-based momentum management policy’s superior performance and robustness. The disturbance estimate was shown to be essential in achieving reliable MPC prediction and bounded momentum management. Furthermore, the proposed MPC policy successfully managed angular momentum growth under maneuvers that exceed the capability of NASA’s state-of-the-art method designed for Solar Cruiser, establishing a larger operational slew envelope. It is shown that the MPC policy proactively unloads angular momentum in preparation for upcoming high-demand slew maneuvers. The framework also proved its efficiency by demonstrating reduced actuator usage through a lower AMT travel distance and optimized RCD usage compared to the benchmark method. This improvement has the potential to enable greater solar sail mission longevity.
Future work on this topic could be the investigation of improving roll disturbance estimates and reducing the number of RCD on-off cycles. Additional work towards the implementation of the proposed method on flight hardware and software will also be pursued to move towards its implementation on future solar sail missions.
@@ Acknowledgments
This material is based upon work supported by NASA under award No. 80NSSC25M7060, as well as a study grant from Chung Cheng Institute of Technology, National Defense University, Taiwan (R.O.C.). The authors would like to thank Mr. Keegan R. Bunker for the valuable discussions on sail shape modeling and providing the SRP model used in this work.
Appendix: Reference Trajectory
For the reference maneuver in this work, the design variables include the initial attitude ${\boldsymbol{\theta}}_{0}$, the predetermined goal attitude ${\boldsymbol{\theta}}_{\text{goal}}$, the maximum allowable slew rate $\dot{{\boldsymbol{\theta}}}_{\text{slew}}\geq\mbf{0}$, and the acceleration time $t_{\text{accel}}$ required to reach the maximum slew rate. The slew maneuver is classified into acceleration, coast, and deceleration phases for both the forward maneuver start at $t_{\text{start}}$ and the return maneuver initiating at $t_{\text{return}}$.
The reference trajectory holds at its initial state ${\boldsymbol{\theta}}_{d}(t)={\boldsymbol{\theta}}_{0}$ and $\dot{{\boldsymbol{\theta}}}_{d}(t)=\mbf{0}$ for $t<t_{\text{start}}$. The system then starts the forward maneuver, slewing toward ${\boldsymbol{\theta}}_{\text{goal}}$, at $t_{\text{start}}$. The constant design variables $\dot{{\boldsymbol{\theta}}}_{\text{slew}}$ and $t_{\text{accel}}$ define a constant angular acceleration magnitude $|\ddot{\theta}_{d,i}|=\dot{\theta}_{\text{slew},i}/t_{\text{accel}}$ for the $i$-th axis. During the acceleration phase where $t-t_{\text{start}}<t_{\text{accel}}$, the angular rate magnitude increases linearly following $|\dot{\theta}_{d,i}|=|\ddot{\theta}_{d,i}|(t-t_{\text{start}})$. Concurrently, the desired attitude evolves quadratically from the initial state as $\theta_{d,i}=\theta_{0,i}\pm\frac{1}{2}|\ddot{\theta}_{d,i}|(t-t_{\text{start}})^{2}$, where the $\pm$ sign depends on the slew direction defined by $\tilde{\theta}_{i}=\theta_{\text{goal},i}-\theta_{0,i}$. Once the maximum slew rate is reached, the system enters a coast phase where $|\dot{\theta}_{d,i}|=\dot{\theta}_{\text{slew},i}$ is maintained. The constant angular velocity is held for $t_{\text{coast}}=2\tilde{\theta}_{i}/\dot{\theta}_{\text{slew},i}-2t_{\text{accel}}$, until the remaining angular distance dictates the start of the deceleration phase. During the deceleration phase, a constant negative acceleration of magnitude $|\ddot{\theta}_{d,i}|$ is applied to smoothly bring the angular rate $\dot{\theta}_{d,i}$ back to zero exactly as the attitude reaches $\theta_{\text{goal},i}$.
The area under the angular velocity versus time plot denotes the total attitude angle change. This integrated area matches the magnitude of slew angular displacement $|\tilde{\theta}_{i}|$. This geometric relationship establishes a strict kinematic constraint that dictates the angular velocity profile. If the slew angle is large enough to complete the full acceleration and deceleration ramps, the standard trapezoidal profile is executed. If the required slew angle is too small, the maximum slew rate cannot be achieved without overshooting the target. Under this condition, the reference trajectory degenerates into a triangular velocity profile. The system accelerates and shortly begins deceleration before $t=t_{\text{start}}+t_{\text{accel}}$ without ever entering a constant velocity coast phase. To satisfy the exact slew angle constraint, the acceleration duration for the degenerated triangular velocity profile is calculated as $t_{\text{accel}\Delta}=\sqrt{|\tilde{\theta}_{i}|/|\ddot{\theta}_{d,i}|}$. The peak angular rate of the slew is thereby reduced to $|\ddot{\theta}_{d,i}|t_{\text{accel}\Delta}$.
Following the completion of the forward maneuver, the system settles and holds at the goal state ${\boldsymbol{\theta}}_{d}={\boldsymbol{\theta}}_{\text{goal}}$, ${\boldsymbol{\omega}}_{d}=\mbf{0}$. This reference orientation is maintained until the specified return time $t_{\text{return}}$ triggers an identical but reversed kinematic sequence to drive the attitude from ${\boldsymbol{\theta}}_{\text{goal}}$ back to the initial state ${\boldsymbol{\theta}}_{0}$ for the interval $t_{\text{complete}}<t<t_{\text{return}}$, where $t_{\text{complete}}=t_{\text{start}}+|\tilde{\theta}_{i}|/\omega_{\text{slew},i}+t_{\text{accel}}$ when the slew angle is large enough to execute the standard trapezoidal profile, and $t_{\text{complete}}=t_{\text{start}}+2t_{\text{accel}\Delta}$ when the slew angle is insufficient and the trajectory degenerates into a triangular profile. This desired slew maneuver trajectory is formulated as a function that calculates the associated reference ${\boldsymbol{\theta}}_{d}(t)$, $\dot{{\boldsymbol{\theta}}}_{d}(t)$, and $\ddot{{\boldsymbol{\theta}}}_{d}(t)$ at any given time $t$. In the RW PID control law, the desired angular momentum is calculated as ${\boldsymbol{\omega}}_{d}(t)=\mbf{S}({\boldsymbol{\theta}}_{d}(t))\dot{{\boldsymbol{\theta}}}_{d}(t)$, where the mapping matrix of a 3-2-1 Euler angle sequence is defined as
$\mbf{S}({\boldsymbol{\theta}})=\begin{bmatrix}1&0&-\sin(\theta_{2})\\ 0&\cos(\theta_{1})&\sin(\theta_{1})\cos(\theta_{2})\\ 0&-\sin(\theta_{1})&\cos(\theta_{1})\cos(\theta_{2})\end{bmatrix}$.